OpenStudy (anonymous):

x^2+18x+4=0 Solve for x, simplify answers using radicals as needed, express complex numbers in terms of i.

OpenStudy (anonymous):

delta = 18 ^2 - 4*4 = 308 x = (-18 + or - sqrt(308)) / 2

OpenStudy (anonymous):

it's not right... is that negative then sqrt?

OpenStudy (anonymous):

why dont u complete the square or use quadratic formula??

OpenStudy (anonymous):

who... me?? That's why I'm posting here... because I don't know these kinds of problems well yet. Need some assistance...

OpenStudy (anonymous):

ok. to use the quad formula simplify -18+or-the sqrt of ((18^2-(4)(1)(4)). get that ans then divide by 2

OpenStudy (anonymous):

x^2 +18x +4 = 0 sqrt(18^2 -4(1)(4) ) = sqrt( 324 -16) =sqrt(308) = sqrt(4*77) = 2sqrt(77) x =( -18 +- 2sqrt(77)) / 2 x = -9 +/- [sqrt(77)]

OpenStudy (anonymous):